# Attributed-graphs kernel implementation using local detuning of neutral-atoms Rydberg Hamiltonian

### v1

@@ Abstract
We extend the quantum-feature kernel framework, which relies on measurements of graph-dependent observables, along three directions. First, leveraging neutral-atom quantum processing units (QPUs), we introduce a scheme that incorporates attributed graphs by embedding edge features into atomic positions and node features into local detuning fields of a Rydberg Hamiltonian. We demonstrate both theoretically and empirically that local detuning enhances kernel expressiveness. Second, in addition to the existing quantum evolution kernel (QEK), which uses global observables, we propose the generalized-distance quantum-correlation (GDQC) kernel, based on local observables. While the two kernels show comparable performance, we show that GDQC can achieve higher expressiveness. Third, instead of restricting to observables at single time steps, we combine information from multiple stages of the quantum evolution via pooling operations. Using extensive simulations on two molecular benchmark datasets, MUTAG and PTC_FM, we find: (a) QEK and GDQC perform competitively with leading classical algorithms; and (b) pooling further improves performance, enabling quantum-feature kernels to surpass classical baselines. These results show that node-feature embedding and kernel designs based on local observables advance quantum-enhanced graph machine learning on neutral-atom devices.
@@ I Introduction
More recently, interest has surged in mapping graphs to quantum systems and computing similarities directly from physical measurements. Early work framed the evolution of a continuous-time quantum walk as a positive-definite kernel on graphs [9, 10], and subsequent extensions employed discrete-time walks and information-theoretic divergences to enrich the feature space [11]. Modern quantum platforms now encode each graph as a parameterized quantum state and estimate kernel entries via repeated state-overlap measurements, a strategy validated experimentally on a 10-qubit nuclear-magnetic-resonance processor [12]. Complementing these hardware demonstrations, industry toolkits such as Pasqal’s open-source Quantum Evolution Kernel (QEK) [13, 14] make quantum-enhanced graph machine-learning pipelines accessible on neutral-atom devices [15]. Together, these developments suggest that quantum hardware may soon deliver graph kernels with higher expressive power and lower sample complexity than their classical counterparts, reinforcing the motivation for the methods introduced here. Despite this notable progress the graph-kernel literature from the perspective of quantum computing remains strongly skewed towards unattributed graphs. Kernels that focus on attributed graphs with discrete labels attached to vertices and edges have been studied comparatively few [16].
Quantum feature kernels such as used in the QEK algorithm represent a particular class of quantum kernels within which the quantum features are derived from measured observables. Herein, a graph $\mathcal{G}$ is embedded into the system Hamiltonian $\hat{H}^{\mathcal{G}}$ such that a time evolution imposes the graph information onto the quantum state $\left|\psi_{\mathcal{G}}\left(t\right)\right\rangle$, and eventually onto any measured observable $O_{\alpha}^{\mathcal{G}}\left(t\right)=\left\langle\psi_{\mathcal{G}}\left(t\right)\right|\hat{O}_{\alpha}\left|\psi_{\mathcal{G}}\left(t\right)\right\rangle$. For the purpose of a quantum feature kernel, multiple observables can be considered ($\alpha=1,\dots,N_{M}$) and measurements can be taken at multiple time steps ($t_{1},\dots,t_{N_{t}}$). In combination, they provide the $\left(N_{M}N_{t}\right)$-dimensional quantum-feature vector $\bm{O}^{\mathcal{G}}$ which encodes the graph information in a non-trivial way. This allows to generically define a quantum feature kernel between two graphs of a given dataset, $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$, as
@@
In this work, we expand on the attributed-graphs toolset by introducing an implementation scheme that allows quantum feature kernels (1) to deal with node attributes. Specifically, we use the local detuning of a neutral-atom Rydberg Hamiltonian to encode node features. We demonstrate that corresponding features kernels (1) are more expressive than without encoded node features, and we present extensive numerical comparison between both scenarios on two well-known benchmark datasets, MUTAG and PTC_FM. Furthermore, this work includes a study of different kernel functions $F_{\kappa}$, whereby two conceptually different implementations are tested, namely QEK and the Generalized-Distance Quantum-Correlation (GDQC) kernel. In contrast to the established QEK kernel which exploits a global observable, the proposed GDQC kernel draws from a local observable, the correlation matrix and is inspired by the classical generalized-distance Weisfeiler-and-Leman (GD-WL) kernel [6]. Lastly, we add one more layer of enhanced expressive power to the project: pooling. Specifically, we combine kernels computed by observables measured at multiple time steps by means of pooling operations. Hereby, two different pooling schemes are tested, sum pooling and product pooling. Having trained a support vector machine (SVM) using all these different kernel implementations, the results of the extensive numerical tests are presented in section IV, and discussed. We observe a trend that the kernels using node features generally yield better SVM predictions compared to the node-agnostic ones. Among the quantum-feature kernels, QEK and GDQC, we observe comparative performances which are competitive with the classical baselines. However, we find that pooling can enhance their performance and enable them to outperform the classical algorithms.
@@
This paper is organized as follows: In Sec. II we detail all aspects of the graph-feature extraction. We introduce the neutral-atom Rydberg Hamiltonian (II.1), introduce the attributed-graphs embedding scheme (II.2), and present all details concerning the emulation (II.3). In Sec. III we discuss classical kernel methods used for the baselines before we introduce the quantum kernel functions $F_{\kappa}$ used in this work. We revisit the QEK kernel (III.1), introduce the proposed GDQC kernel (III.2), before we define the two pooling schemes in part III.3. In Sec. IV, the numerical results are presented and discussed. Finally, we provide a conclusion in Sec. V. In App. A we rigorously demonstrate that the kernels with node-feature embedding are more expressive than the ones without. Apps. B and C prove, respectively, that the GDQC kernel satisfies all kernel properties and that it is as expressive as the GD-WL kernel. In App. D we prove that the inclusion of different time steps can only increase the kernel expressiveness.
@@ II Quantum feature extraction
In this section we demonstrate how the quantum features $\bm{O}^{\mathcal{G}}$ [cf. Eq. (1)] are extracted using a neutral-atom QPU. After introducing the Rydberg Hamiltonian, we discuss the graph embedding scheme, and explain the details of the emulation used to extract the graph-dependent observables.
@@ II.1 Rydberg/ Ising Hamiltonian
where the projection operator onto a fixed-excitation subspace is defined as $\hat{P}_{k}=\sum_{\bm{b}}\delta_{k,\sum_{i}b_{i}}\,\left|\bm{b}\right\rangle\left\langle\bm{b}\right|$ with the Kronecker delta $\delta_{k,k^{\prime}}$ and $k=0,1,\dots,N$. All observables (6) are diagonal in the measurement basis, and thus, they can be easily sampled after each run of the neutral-atom QPU.
@@ II.2 Attributed-graphs embedding
A common strategy to embed a graph onto the Hamiltonian (2) is by choosing the atom’s positions such that connected atoms have a significant coupling strength $J_{ij}$ whereas unconnected atoms are coupled with a negligible strength. Mathematically, we introduce the graph-dependent coupling matrix as
@@
where we defined the nearest-neighbor coupling strength $J_{\mathrm{NN}}=C_{6}/r_{\mathrm{NN}}^{6}$ with $r_{\mathrm{NN}}$ being the nearest-neighbor distance. Throughout this study, we fix the nearest-neighbor distance to $r_{\mathrm{NN}}=5\,\mathrm{\mu m}$ such that the nearest-neighbor coupling becomes $J_{\mathrm{NN}}/\hbar\approx 55.4\,\mathrm{MHz}$. The choice (7) is very appealing for the graph embedding. Yet, due to the Rydberg blockade mechanism [21, 22] only a certain class of graphs can be embedded via this scheme. These are the so-called unit-disk graphs wherein an edge between two nodes only exists if the distance between them is smaller than a fixed value [23]. In general, the task of finding a unit-disk graph embedding is a NP-complete problem [24]. Conveniently, the molecules in the dataset just like the atoms in the QPU interact through physical forces which naturally makes them unit-ball graphs. The only problem is the dimension. Current neutral-atom QPU only allow for two-dimensional registers. Therefore, in our study we only include molecules that admit a two-dimensional embedding reducing the size of the datasets to $171$ (MUTAG) and $312$ (PTC_FM) molecules. We denote the reduced datasets with an asterisk . The algorithm used to compute the layouts is an extension of the one described in Ref. [14] (appendix E therein).
Figure 1: Emulation results for the molecule in (a) consisting of five carbon (blue) and two oxygen (red) atoms. (b)-(d) Time evolution of the three observables introduced in Eq. (6), the excitation probability $P_{k}$, the excitation number $n_{i}=\langle\hat{n}_{i}\rangle$ and the pairwise correlations $C_{ij}=\langle\hat{n}_{i}\hat{n}_{j}\rangle$. The solid (dashed) lines correspond to evolution subjected to the global-detuning (local-detuning) Hamiltonian, see Eq. (9) (Eq. (10)). For the given molecule and parameter values one observes differences between the two detuning cases starting at the latest after $t\gtrsim 250\,\mathrm{ns}$. These differences can be learned by the machine learning models designed in trained in Secs. III and IV. The molecule in (a) belongs to the PTC_FM dataset, and it has the SMILES representation "CC1(C)COC1=O".
@@
$\delta_{i}^{\mathcal{G}}$ $=\delta_{0}\,m_{C}\big{/}m_{i}\quad\leq\delta_{0},$ (8)
where $\delta_{0}$ is a fixed offset value and $m_{C}=12\,\mathrm{u}$ is the mass of a carbon atom, the lightest atom in the datasets (Note that hydrogen atoms are omitted in the graphs.). The atomic masses are obtained from PubChem Periodic Table webpage [25]. This allows us to study the graph-embedded Hamiltonian (2), once with the inclusion of the node-feature embedding (8), and once without. We define the respective Hamiltonians as the (global) Hamiltonian
@@
which is subjected to a globally fixed detuning value $\bm{\delta}_{\mathrm{glob}}=\left(\delta_{0},\dots,\delta_{0}\right)$, and as the (local) Hamiltonian
@@
$h_{i}^{z}$ $\approx\frac{1}{4\hbar}J_{\mathrm{NN}}\left(-N_{i}^{\mathrm{deg}}+\frac{2\hbar\delta_{0}}{J_{\mathrm{NN}}}\,\frac{m_{C}}{m_{i}}\right).$ (11)
Here, we note two things. First, in order for the local-detuning to be noticeable, one has to choose the offset value such that $2\hbar\delta_{0}/J_{\mathrm{NN}}=\mathcal{O}\left(1\right)$. Second, we see that the offset value conveniently encodes the degree of each node as a default piece of information. We would not want to overshadow this embedded node information, but instead add a little value on top according to the node feature, i.e. according to the mass of atom $i$. For this reason, we consider an appropriate choice for the fixed detuning value as $2\hbar\delta_{0}/J_{\mathrm{NN}}\lesssim 0.3$. The value we choose in this work is $\delta_{0}=\pi\,\mathrm{rad}\,/\mathrm{\mu s}$, such that $2\hbar\delta_{0}/J_{\mathrm{NN}}\approx 0.1$.
@@ II.3 Emulation
With all the specifications provided we run two emulations for every graph in the reduced datasets MUTAG and PTC_FM: Once where the time evolution is governed by the local-detuning Hamiltonian (10) and once governed by the global-detuning counterpart (9). At multiple time steps, we sample the three observables defined in Eq. (6), the probability for a fixed number of excitations $P_{k}$, the excitation number $n_{i}$ and the correlation $C_{ij}$, with a number of shots $N_{\mathrm{shots}}=1000$. Their values form the set of feature vectors $\bm{O}^{\mathcal{G}_{\mu}}$, see Eq. (1), which are used in the machine learning tasks in Secs. IV and III. The emulation is computed using two open-source libraries: Pulser [26] and EMU-MPS [27]. Pulser is a Python toolkit for neutral-atom quantum computing that allows to program experiments directly at the pulse level. The package EMU-MPS provides an efficient, state-of-the-art emulation backend that emulates pulse-level dynamics of neutral-atom arrays by storing the state as a matrix product state.
In Fig. 1 we demonstrate the time evolution by presenting emulation results for a given molecule [panel (a)] of the PTC_FM dataset. The three panels (a)-(c) show respectively, the excitation probability $P_{k}$, the excitation number $n_{i}$ and the correlation matrix elements $C_{ij}=\langle\hat{n}_{i}\hat{n}_{j}\rangle$. To highlight the effect of the local detuning we compare the local-detuning results (dashed lines) with the global-detuning ones (solid lines). In all three observables one observes differences between the two cases emerging at the latest after $t\gtrsim 250\,\mathrm{ns}$. Note that the presented molecule consists of five carbon and two oxygen molecules. The given implementation scheme is such that the global-detuning Hamiltonian (9) effectively assumes that all atoms are carbon atoms ($m_{\mathrm{C}}=12\,\mathrm{u}$). Thus, the oxygen nodes (0 and $6$) are particularly interesting. Among the two excitation numbers $n_{0}$ and $n_{6}$, only the node $n_{0}$ with node degree two shows a significant deviation. During the research phase we have also tested a larger detuning value of $\delta_{0}=2\pi\,\mathrm{rad}\,/\mathrm{\mu s}$ which also enforces a strong deviation for $n_{6}$. Across the entire datasets however, this has only a minor effect as the prediction tasks in Sec. IV yield similar results. Note also that atoms heavier than oxygen ($m_{\mathrm{O}}=16\,\mathrm{u}$) are significantly stronger influenced by the local detuning.

### v2

@@ Abstract
We extend the quantum-feature kernel framework, which relies on measurements of graph-dependent observables, along three directions. First, leveraging neutral-atom quantum processing units (QPUs), we introduce a scheme that incorporates attributed graphs by embedding edge features into atomic positions and node features into local detuning fields of a Rydberg Hamiltonian. We demonstrate both theoretically and empirically that local detuning enhances kernel expressiveness. Second, in addition to the existing quantum evolution kernel ($\mathrm{QEK}$), which uses global observables, we propose the generalized-distance quantum-correlation ($\mathrm{GDQC}$) kernel, based on local observables. While the two kernels show comparable performance, we show that $\mathrm{GDQC}$ can achieve higher expressiveness. Third, instead of restricting to observables at single time steps, we combine information from multiple stages of the quantum evolution via pooling operations. Using extensive simulations on two molecular benchmark datasets, MUTAG and PTC_FM, we find: (a) $\mathrm{QEK}$ and $\mathrm{GDQC}$ perform competitively with leading classical algorithms; and (b) pooling further improves performance, enabling quantum-feature kernels to surpass classical baselines. These results show that node-feature embedding and kernel designs based on local observables advance quantum-enhanced graph machine learning on neutral-atom devices.
@@ I Introduction
More recently, interest has surged in mapping graphs to quantum systems and computing similarities directly from physical measurements. Early work framed the evolution of a continuous-time quantum walk as a positive-definite kernel on graphs [41, 51], and subsequent extensions employed discrete-time walks and information-theoretic divergences to enrich the feature space [6]. Modern quantum platforms now encode each graph as a parameterised quantum state and estimate kernel entries via repeated state-overlap measurements [30], a strategy validated experimentally on a 10-qubit nuclear-magnetic-resonance processor [42]. Complementing these hardware demonstrations, industry toolkits such as Pasqal’s open-source Quantum Evolution Kernel ($\mathrm{QEK}$[24, 2] compute graph kernels from expectation values [32] of graph-dependent observables measured on neutral-atom devices [35]. Together, these developments suggest that quantum hardware may soon deliver graph kernels with higher expressive power and lower sample complexity than their classical counterparts, reinforcing the motivation for the methods introduced here. Despite this notable progress, the graph-kernel literature from the perspective of quantum computing remains strongly skewed towards unattributed graphs. Kernels that focus on attributed graphs with discrete labels attached to vertices and edges have been comparatively few [5].
Quantum feature kernels such as used in the $\mathrm{QEK}$ algorithm represent a particular class of quantum kernels within which the quantum features are derived from measured observables. Herein, a graph $\mathcal{G}$ is embedded into the system Hamiltonian $\hat{H}^{\mathcal{G}}$ such that a time evolution imposes the graph information onto the quantum state $\left|\psi_{\mathcal{G}}\left(t\right)\right\rangle$, and eventually onto any measured observable $O_{\alpha}^{\mathcal{G}}\left(t\right)=\left\langle\psi_{\mathcal{G}}\left(t\right)\right|\hat{O}_{\alpha}\left|\psi_{\mathcal{G}}\left(t\right)\right\rangle$. For the purpose of a quantum feature kernel, multiple observables can be considered ($\alpha=1,\dots,N_{M}$) and measurements can be taken at multiple time steps ($t_{1},\dots,t_{N_{t}}$). In combination, they provide the $\left(N_{M}N_{t}\right)$-dimensional quantum-feature vector $\bm{O}^{\mathcal{G}}$ which encodes the graph information in a non-trivial way. This allows to generically define a quantum feature kernel between two graphs of a given dataset, $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$, as
@@
In this work, we expand on the set of attributed-graphs tools by introducing an implementation scheme that allows quantum feature kernels (1) to deal with node attributes. Specifically, we used the local detuning of a neutral-atom Rydberg Hamiltonian to encode node features. We demonstrate that the corresponding feature kernels (1) are more expressive than without encoded node features, and we present an extensive numerical comparison between both scenarios on two well-known benchmark datasets, MUTAG and PTC_FM. Furthermore, this work includes a study of different kernel functions $F_{\kappa}$, in which two conceptually different implementations are tested, namely $\mathrm{QEK}$ and the Generalized-Distance Quantum-Correlation ($\mathrm{GDQC}$) kernel. In contrast to the established $\mathrm{QEK}$ kernel which exploits a global observable, the proposed $\mathrm{GDQC}$ kernel draws from a local observable, the correlation matrix, and is inspired by the classical generalized-distance Weisfeiler-and-Leman (GD-WL) kernel [47]. Lastly, we add one more layer of enhanced expressive power to the project: pooling. Specifically, we combine kernels computed by observables measured at multiple time steps by means of pooling operations. Hereby, two different pooling schemes are tested, sum pooling and product pooling. Having trained a support vector machine (SVM) using all these different kernel implementations, the results of the extensive numerical tests are presented in section IV, and discussed. We observe a trend that the kernels using node features generally yield better SVM predictions compared to the node-agnostic ones. Among the quantum-feature kernels, $\mathrm{QEK}$ and $\mathrm{GDQC}$, we observe competitive performances with the classical baselines. However, we find that pooling can enhance their performance and enable them to outperform the classical algorithms.
@@
This paper is organized as follows: In Sec. II we detail all aspects of the graph-feature extraction. We introduce the neutral-atom Rydberg Hamiltonian (II.1), introduce the attributed-graphs embedding scheme (II.2), present all details concerning the emulation (II.3), and show analytical small-time approximations useful for performance interpretation (II.4). In Sec. III we discuss the implemented kernel methods used in this work. First, we revisit classical kernel methods (III.1), before we introduce the $\mathrm{QEK}$ kernel (III.2), and the proposed $\mathrm{GDQC}$ kernel (III.3), and lastly in Subsec. III.4, we define the two implemented pooling schemes. In Sec. IV, the numerical results are presented and discussed. Finally, we provide a conclusion in Sec. V. In App. A we derive the analytical small-time approximations via time-dependent perturbation theory. App. B provides practical details on the classical-kernel implementations. In App. C we rigorously demonstrate that the kernels with node-feature embedding are more expressive than the ones without. Apps. D and E prove, respectively, that the $\mathrm{GDQC}$ kernel satisfies all kernel properties and that it is at least as expressive as the GD-WL kernel. In App. F we prove that the inclusion of different time steps can only increase the kernel expressiveness.
@@ II Quantum feature extraction
In this section we demonstrate how the quantum features $\bm{O}^{\mathcal{G}}$ [cf. Eq. (1)] are extracted using a neutral-atom QPU. After introducing the Rydberg Hamiltonian (II.1), we discuss the graph embedding scheme (II.2), and explain the details of the emulation used to extract the graph-dependent observables (II.3). Finally, we present small-time analytical approximations useful for later results‘ interpretation (II.4).
@@ II.1 Rydberg / Ising Hamiltonian
where the projection operator onto a fixed-excitation subspace is defined as $\hat{P}_{k}=\sum_{\bm{b}}\delta_{k,\sum_{i=1}^{N}b_{i}}\,\left|\bm{b}\right\rangle\left\langle\bm{b}\right|$ with the Kronecker delta $\delta_{k,k^{\prime}}$ and $k=0,1,\dots,N$. Note that since $\hat{n}_{i}^{2}=\hat{n}_{i}$, the diagonal elements of the correlation matrix $C_{ii}\left(t\right)=n_{i}\left(t\right)$ are identical to the excitation numbers $n_{i}\left(t\right)$, also referred to as the site occupations. All observables (6) are diagonal in the measurement basis, and thus, they can be easily sampled after each run of the neutral-atom QPU.
The Rydberg blockade mechanism [11, 1] provides a useful estimate of the length scale over which Rydberg-Rydberg interactions suppress simultaneous excitation of atoms. Equating the van-der-Waals interaction shift to the drive energy scale, $C_{6}/R_{\mathrm{b}}^{6}=\hbar\Omega$, defines the blockade radius $R_{\mathrm{b}}=(C_{6}/\hbar\Omega)^{1/6}$. For interatomic separations $r_{ij}\lesssim R_{\mathrm{b}}$, excitation of the doubly excited state is energetically suppressed (shifted out of resonance). In our setting, we choose $R_{\mathrm{b}}=7.19\,\mathrm{\mu m}$ with a nearest-neighbor distance $r_{\mathrm{NN}}=5.01\,\mathrm{\mu m}$ (see below) such that $r_{\mathrm{NN}}<R_{\mathrm{b}}<2r_{\mathrm{NN}}$, which blockades nearest neighbors while leaving neighbors at further distances largely unblocked.
@@ II.2 Attributed-graphs embedding
A common strategy to embed a graph onto the Hamiltonian (2) is by choosing the atom‘s positions such that connected atoms have a significant coupling strength $J_{ij}$ whereas unconnected atoms are coupled with a negligible strength. Mathematically, we introduce the graph-dependent coupling matrix as
@@
where we defined the nearest-neighbor coupling strength $J_{\mathrm{NN}}=C_{6}/r_{\mathrm{NN}}^{6}$ with $r_{\mathrm{NN}}$ being the nearest-neighbor distance. Throughout this study, we fix the nearest-neighbor distance to $r_{\mathrm{NN}}=5.01\,\mathrm{\mu m}$ such that the nearest-neighbor coupling becomes $J_{\mathrm{NN}}/\hbar\approx 54.75\,\mathrm{rad}/\mu s$. The choice in Eq. (7) is appealing for graph embedding, although it restricts this scheme to a specific class of graphs, namely unit-disk graphs. Here, an edge between two nodes exists only if their distance is below a fixed threshold distance [12]. In general, the task of finding a unit-disk graph embedding is a NP-complete problem [28]. Conveniently, the molecules in the dataset just like the atoms in the QPU interact through physical forces which naturally makes them unit-ball graphs. The only problem is the dimension. Current neutral-atom QPU only allow for two-dimensional registers. Therefore, in our study we only include molecules that admit a two-dimensional embedding reducing the size of PTC_FM dataset to $332$ molecules, corresponding to $95\%$ of the original dataset. The MUTAG dataset was kept intact, since we obtained a two-dimensional representation of all 188 constituent molecules. We denote the reduced PTC_FM dataset as PTC_FM. The algorithm used to compute the layouts is an extension of the one described in Ref. [2] (appendix E therein), i.e. a force-based algorithm [16, 26].
Figure 1: Emulation results for the molecule in panel (a). Panels (b)-(d) display the time evolution of the three observables defined in Eq. (6): the excitation probability $P_{k}$, the site occupations $n_{i}=\langle\hat{n}_{i}\rangle$, and the pairwise correlations $C_{i,j}=\langle\hat{n}_{i}\hat{n}_{j}\rangle$. Solid (dashed) curves correspond to evolution under the global-detuning (local-detuning) Hamiltonian, with detuning scale set to $\delta_{0}=2\pi\,\mathrm{rad}\,/\mathrm{\mu s}$; see Eqs. (9) and (10). For comparison, the inset in (d) shows results for a smaller detuning scale, $\delta_{0}=\pi\,\mathrm{rad}\,/\mathrm{\mu s}$. For the molecule and parameters shown, differences between global and local detuning become apparent by $t\gtrsim 250\,\mathrm{ns}$. The molecule in (a) is taken from the PTC_FM dataset and has SMILES string "CC1(C)COC1=O". The gray dotted vertical line marks $t=0.4\,\mathrm{\mu s}$, which is the time step used for the visualization in Fig. 2.
@@
$\delta_{i}^{\mathcal{G}}$ $=\delta_{0}\left(1-\frac{m_{C}}{m_{i}}\right)$ (8)
where $\delta_{0}$ is a fixed offset value and $m_{C}=12\,\mathrm{u}$ is the mass of a carbon atom, which is the lightest element present in our datasets (hydrogen atoms are omitted in the graphs.). The atomic masses are obtained from PubChem Periodic Table webpage [39]. The local detuning choice (8) is motivated as follows. In earlier $\mathrm{QEK}$ work [2] the global detuning was set to zero. To ensure that the local detuning provides additional expressivity beyond this global baseline, we design Eq. (8) such that the detuning remains zero for the most abundant element in the datasets. This is the case for carbon, which accounts for approximately $70\%$ of all atoms. Consequently, only the remaining atoms receive a nonzero local detuning, i.e., $\delta_{i}^{\mathcal{G}}>0$ whenever $m_{i}>m_{C}$. The scheme in Eq. (8) enables us to study the graph-embedded Hamiltonian in Eq. (2) both with and without the node-feature (local detuning) embedding, allowing for a direct, controlled comparison between the two settings.
We define the respective Hamiltonians as the (global) Hamiltonian
@@
which is subjected to a globally fixed detuning value $\bm{\delta}_{\mathrm{glob}}=\left(0,\dots,0\right)$, and as the (local) Hamiltonian
@@
$h_{i}^{z}$ $\approx\frac{1}{4\hbar}J_{\mathrm{NN}}\left[-N_{i}^{\mathrm{deg}}+\frac{2\hbar\delta_{0}}{J_{\mathrm{NN}}}\,\left(1-\frac{m_{C}}{m_{i}}\right)\right].$ (11)
Here, we note two things. First, in order for the local-detuning to be noticeable, one has to choose the offset value such that $2\hbar\delta_{0}/J_{\mathrm{NN}}\sim\mathcal{O}\left(1\right)$. Second, we see that the offset value conveniently encodes the degree of each node as a default piece of information. We would not want to overshadow this embedded node information, but instead add a little value on top according to the node feature, i.e. according to the mass of atom $i$. For this reason, we consider an appropriate choice for the fixed detuning value as $2\hbar\delta_{0}/J_{\mathrm{NN}}\lesssim 0.3$. The value we choose in this work is $\delta_{0}=2\pi\,\mathrm{rad}\,/\mathrm{\mu s}$, such that $2\hbar\delta_{0}/J_{\mathrm{NN}}\approx 0.23$.
@@ II.3 Emulation
With all the specifications provided we run two emulations for every graph in the datasets MUTAG and PTC_FM: Once where the time evolution is governed by the local-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10) and once governed by the global-detuning counterpart $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9). At multiple time steps, we sample the three observables defined in Eq. (6), the probability for a fixed number of excitations $P_{k}$, the excitation number $n_{i}$ and the correlation $C_{ij}$, with a number of shots $N_{\mathrm{shots}}=1000$. Their values form the set of feature vectors $\bm{O}^{\mathcal{G}_{\mu}}$, see Eq. (1), which are used in the machine learning tasks in Secs. IV and III. The emulation is computed using two open-source libraries: Pulser [50] and EMU-MPS [7]. Pulser is a Python toolkit for neutral-atom quantum computing that allows to program experiments directly at the pulse level. The package EMU-MPS provides an efficient, state-of-the-art emulation backend that emulates pulse-level dynamics of neutral-atom arrays by storing the state as a matrix product state.
In Fig. 1, we illustrate the time evolution by showing emulation results for a representative molecule from the PTC_FM dataset [panel (a)]. Panels (b)-(d) report, respectively, the excitation probability $P_{k}$, the site occupations $n_{i}$, and the correlation matrix elements $C_{i,j}=\langle\hat{n}_{i}\hat{n}_{j}\rangle$, cf. Eq. (6). The main purpose of this figure is to highlight the effect of the local detuning. To this end, we compare results obtained with global detuning (solid lines) to those obtained with local detuning (dashed lines) at $\delta_{0}=2\pi\,\mathrm{rad}\,/\mathrm{\mu s}$, cf. Eq. (8). Across all three observables, deviations between the two settings become visible after $t\gtrsim 250\,\mathrm{ns}$. The molecule shown in panel (a) contains five carbon atoms and two oxygen atoms. Under the global-detuning Hamiltonian, Eq. (9), all atoms effectively behave as carbon atoms ($m_{\mathrm{C}}=12\,\mathrm{u}$). Consequently, the oxygen nodes (0 and $6$) are particularly informative. Consistent with this, Fig. 1(c) shows that the carbon site $n_{2}$, which is distant from the oxygen atoms, is largely unaffected, whereas the oxygen sites ($n_{0}$ and $n_{6}$) and nearby carbon sites (e.g., $n_{5}$) exhibit pronounced differences between global and local detuning. Finally, to emphasize the role of the detuning scale $\delta_{0}$, the inset of Fig. 1(d) shows the local-detuning correlation matrix results for a smaller value, $\delta_{0}=\pi\,\mathrm{rad}\,/\mathrm{\mu s}$. In this regime, global and local-detuning observables differ only marginally, so the resulting signal is likely too weak for a machine learning model to leverage. As mentioned before, in this work we set the scale to $\delta_{0}=2\pi\,\mathrm{rad}\,/\mathrm{\mu s}$.
II.4 Time-dependent perturbation theory
In this part we conduct an early-time expansion for the relevant observables (6). Specifically, we treat the transverse-field part of the Hamiltonian (2) as a perturbation, i.e. we assume $\Omega t\ll 1$. The resulting early-time approximations allow us to gain useful insights into the performance of the quantum kernels introduced in Sec. III. Details on the derivation of the perturbation theory can be found in App. A.
Neglecting all terms of order $\mathcal{O}\left((\Omega t)^{4}\right)$ and higher, the leading observable contributions are given as
$P_{0}\left(t\right)$ $=1-\frac{\Omega^{2}t^{2}}{4}\sum_{i=1}^{N}\mathrm{sinc}^{2}\!\left(\delta_{i}t/2\right)+\mathcal{O}\left((\Omega t)^{4}\right),$ (12)
$P_{1}\left(t\right)$ $=\frac{\Omega^{2}t^{2}}{4}\sum_{i=1}^{N}\mathrm{sinc}^{2}\!\left(\delta_{i}t/2\right)+\mathcal{O}\left((\Omega t)^{4}\right),$ (13)
$n_{i}\left(t\right)$ $=\frac{\Omega^{2}t^{2}}{4}\mathrm{sinc}^{2}\!\left(\delta_{i}t/2\right)+\mathcal{O}\left((\Omega t)^{4}\right),$ (14)
where $\mathrm{sinc}\left(x\right)=\sin\left(x\right)/x$. As shown in App. A the excitation probability scales as $P_{k}\left(t\right)=\mathcal{O}\left((\Omega t)^{2k}\right)$, such that $P_{k\geq 2}\left(t\right)$ can be neglected within the current approximation. The parabolic expressions for $P_{0}\left(t\right)$ (12) and $P_{1}\left(t\right)$ (13) are consistent with Fig. 1(b). The diagonal elements of the correlation matrix $C_{ii}\left(t\right)=n_{i}\left(t\right)$ are approximated via Eq. (14), whereas the off-diagonal elements scale as $\mathcal{O}\left((\Omega t)^{4}\right)$ consistent with the observed behavior in Fig. 1(d). For early-time dynamics, where $\delta_{i}t\ll 1$, one can further approximate $\mathrm{sinc}\!\left(\delta_{i}t/2\right)=1$, such that all site occupations $n_{i}\left(t\ll\delta_{i}^{-1}\right)$ behave identically, as can be seen in Fig. 1(c).